Determining hours an IV will run:

1.. Order: 1000ml NS IV. Infuse at 50ml/hr. How long will it take for this infusion to go in?

2. Order 500 ml D5W IV. Infuse at 40ml/hr. How many hours will this fluid infuse?

3. Order 750 ml LR IV. Infuse at 75ml/hr. How long will it take this infusion to go in?

4.. Order 2000ml NS IV. Infuse at 200ml/hr. How long will it take this infusion to go in?
Answer #1

Ans) Formula:

ml/hr = Total volume to be administered ÷ time in hrs

1) Order: 1000ml NS IV. Infuse at 50ml/hr.

50 = 1000 ÷ time

Time = 1000 ÷ 50 = 20 hrs.

2) Order 500 ml D5W IV. Infuse at 40ml/hr.

40 = 500 ÷ time

Time= 500 ÷ 40 = 12.5 hrs.

3) Order 750 ml LR IV. Infuse at 75ml/hr.

75 = 750 ÷ time

Time = 750 ÷ 75 = 10 hrs.

4) Order 2000ml NS IV. Infuse at 200ml/hr.

200 = 2000 ÷ time

Time = 2000 ÷ 200 = 10 hrs.